OpenStudy (anonymous):

find the fifth term of the sequence which has a first term of 17 and a common ratio of 3.1

OpenStudy (anonymous):

@jordanloveangel

OpenStudy (jordanloveangel):

sorry i took long time

OpenStudy (jordanloveangel):

A{1} = 17 ; and a geometric ratio o f 3.1 A{2} = 17 * 3.1 A{3} = 17 * (3.1 * 3.1) A{4} = 17 * (3.1 * 3.1 * 3.1); [3.1^3] A{n} = 17 * 3.1^(n-1) A{5} = 17 * 3.1^(4) I get 1569.98... if i did it right

OpenStudy (anonymous):

It's alright. Thank you omg.

OpenStudy (jordanloveangel):

lol np

OpenStudy (anonymous):

That wasn't right, but I can probably get my mom to help. Thanks for attempting to help though.

OpenStudy (jordanloveangel):

np sorry i tryed

OpenStudy (anonymous):

It's alright.

OpenStudy (jordanloveangel):

@Catlover5925

OpenStudy (catlover5925):

ok so multiply # by 3 17 * 3 is 51 so it is a 3:1 ration so it would be 17, 51, 153, 459, 1377.....i think this is correct

OpenStudy (catlover5925):

does this make sense?? @jxss.xca

OpenStudy (anonymous):

Yeah. Thanks.